OpenStudy (anonymous):

i need help with the elimination method. My two equations are 2x+3y=6 and x+2y=5. What do i do?

OpenStudy (anonymous):

try to write in this form y=mx+b

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

no multiply the second equation by 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

actually by -2 but she is guiding you correctly

OpenStudy (anonymous):

then subtract it from the first equation

OpenStudy (anonymous):

2x+4y=10

OpenStudy (anonymous):

subtract it from 2x+3y=6

OpenStudy (anonymous):

we get y=4

OpenStudy (anonymous):

substitute the value of y in the first equation

OpenStudy (anonymous):

we get 2x+12=6

OpenStudy (anonymous):

2x=-6 x=-3